OpenStudy (anonymous):

Find two consecutive whole numbers such that the sum of their squares is 13

OpenStudy (anonymous):

Putting this into algebra, \[x^2 +(x+1)^2 = 13\], where x is the unknown Solve for x.

OpenStudy (anonymous):

I still can't find the two values..

OpenStudy (anonymous):

hmm? what was you last step?

OpenStudy (anonymous):

I got to 2x^2-13+1=0

OpenStudy (anonymous):

Ah. I see. Note that: \[(x+1)^2 = x^2 + 2x +1\] instead of \(x^2 + 1\)

OpenStudy (anonymous):

I'm still having trouble finding 2 whole numbers... I got -1 + or - radical 13 What do you get when you solve it?

OpenStudy (anonymous):

...I'm also kinda in a hurry, so if you could just tell me the answer it would really help out a lot

OpenStudy (anonymous):

Alright, expanding the original equation, \[x^2 + x^2+2x+1 =13 \] \[2x^2 +2x-12=0\] \[x^2 +x-6 =0 \] I think you can progress from here.

OpenStudy (anonymous):

x=3 or x=-2(Not accepted) so x=3

OpenStudy (anonymous):

I see now, I forgot the other x^2.. and if only 1 answer is accepted, why do I still have to provide 2 consecutive whole numbers for the question?

OpenStudy (anonymous):

because x=2 then x+1 =3? (note:I made a mistake up there, sorry, lol)

OpenStudy (anonymous):

lol its ok, I figured

OpenStudy (anonymous):

it worked. thanks a lot

OpenStudy (anonymous):

You're welcome :)